Reports 1-1 of 1 Clear search Modify search
MIF (ASC)
hirose.chiaki - 5:53 Saturday 05 November 2022 (22801) Print this report
Checked IY and EY sensors of ASC Yarm

Thanks for the advice, Kenta.

This work was on Thursday and Friday.
I measured sensing matrix again for checking IY and EY sensors of ASC Yarm.
Result is "MATRIX".
(../../users/Commissioning/data/ASC/2022/REFL_WFS/Yarm/Sensmat/TF_ASC_Yarm_REFLWFS_OPLEV_{ITMY,ETMY}_{P,Y}.xml)
The results were similar to https://klog.icrr.u-tokyo.ac.jp/osl/?r=22602 in terms of ratio of results and positive/negative.
This result corrected for the phase rotation due to oplev gain switch.

With the matrix unchanged, I measured PLANT and found no coherence.
(../../users/Commissioning/data/ASC/2022/REFL_WFS/Yarm/PLANT/PLANT_{ITMY,ETMY}_{DSOFT,DHARD}_{PIT,YAW}.xml)
I made ITMY and ETMY sensors in DHARD/DSOFT from the results of re-measuring the sensing matrix.(PIT/YAW_ASC_MATRIX_C.png)
I measured PLANT in the same way, but coherence did not appear as well.
(../../users/Commissioning/data/ASC/2022/REFL_WFS/Yarm/PLANT/PLANT_{ITMY,ETMY}_{DSOFT,DHARD}_{PIT,YAW}_C.xml)
I felt the mirror was not tilted.(ITMY_P.png)
I'm not sure how to solve this problem.
I reverted the matrix parameters.

Images attached to this report
Non-image files attached to this report
Search Help
×

Warning

×